Find the central force necessary to make a particle describe the lemniscate r^2 = a^2 cos 2θ, where a is a constant


1
Expert's answer
2021-02-23T10:05:25-0500

Basic equation of celestial mechanics:

"\dfrac{d^2u}{d\theta^{2}}+u=\dfrac{-f(1/u)}{mh^2u^2}" (1)


Equation of lemniscate,

"r^2=a^2cos2\theta" (2)

Put "r=1\/u" in equation (2), we get

"\dfrac{1}{u^2}=a^2cos2\theta"

"\Rightarrow u=\dfrac{1}{a}\sqrt{sec2\theta}" (3)

Differentiating equation (3) with respect to "\theta" , we get

"\dfrac{du}{d\theta}=\dfrac{1}{a}\times\dfrac{1}{\cancel2\sqrt{sec2\theta}}\times sec2\theta\times tan2\theta\times\cancel2"

"\Rightarrow\dfrac{du}{d\theta}=\dfrac{1}{a}sec2\theta^{1/2}tan2\theta" (4)

Again differentiating the above equation

"\dfrac{d^2u}{d\theta^2}=\dfrac{1}{a}(sec^{1/2}2\theta tan^{2}2\theta+2sec^{5/2}2\theta)" (5)

Adding equation (3) and (5) we get

"\dfrac{d^2u}{d\theta^2}+u=\dfrac{1}{a}[sec^{1/2}2\theta \{tan^22\theta +1\}+2sec^{5/2}2\theta]"

"\Rightarrow\dfrac{d^2u}{d\theta^2}+u=\dfrac{1}{a}[3sec^{5/2}2\theta]"

"=3\times\dfrac{1}{a}\times \dfrac{a^5}{r^5}" from eq. (3) and (2)

"=3a^4r^{-5}" (6)


Substituting value of equation (6) in (1)

"3a^4r^{-5}=\dfrac{-f(r)}{mh^2r^{-2}}"

"\Rightarrow f(r)=\dfrac{-3a^4mh^2}{r^7}"

"\therefore" The force f(r) will make a particle describe lemniscate shaped orbit.

"\dfrac{}{}"